To start an avalanche on a mountain slope, an artillery shell is fired with an initial velocity of 332 m/s at 37.0° above the horizontal. It explodes on the mountainside 28.4 s after firing. What is the x component of the shell where it explodes, relative to its firing point? Answer: Check What is the y component of the shell where it explodes, relative to its firing point? Answer: Check
Answer #1

x component = u cos(theta ) = 332 cos(37) = 265.15 m/s

y component = u sin(theta) -g t = 332 sin(37) - (9.81 x 28.4) = -78.52 m/s

please rate it up thanks :)
Answer #2

SOLUTION :



Shell is fired at an angle of 37º and speed of 332 m/s. 

upward. It travels for 28.4 seconds and explodes.


So, 


horizontal speed  of shell , Vh = 332 cos (37) =  265.1470 m/s


vertical speed  of shell  , Vv = 332 sin (37) = 199.8026  m/s


Coordinates of explosion point relative to firing point(0, 0) :


X - component of shell at explosion point 

= Horizontal distance moved by shell 

= Horizontal speed * time 

= 265.1470 * 28.4

 = 7530.17 meters  (ANSWER)


Y-component of shell at explosion point 

= vertical distance travelled by shell 

= u t - 1/2 g t^2   ;  (u = Vv)

= (199.8026)(28.4) - 1/2 * 9.8 * (28.4)^2 

= 1722.25  meters. (ANSWER).



answered by: Tulsiram Garg